Need answers for the highlighted cells in question 1 and 2 (Formula's)JPL, Inc. has provided its sales and expense data for the most recent period. The Controller has asked you prepare a spreadsh? X XS-a. FILE HOME CVP Analysis - Excel FORMULAS DATA - Sign In INSERT PAGE LAYOUT REVIEW VIEW Calibri Paste BIU. - - A Cell? - CVP Analysis - Excel FORMULAS DATA FILE HOME INSERT PAGE LAYOUT REVIEW VIEW Sign In Cells Editing points Paste B I U - 2-Jay Corporation has provided data from a two-year period to aid in planning. The Controller has asked you to prepare a contri? X XS-a, FILE HOME High-Low Method; Contribution Format Income Statement - Excel PAGE LAYOUT FORMULAS DATA REVIEW VIEW - Sig? X - Sign In XS-a, High-Low Method; Contribution Format Income Statement - Excel FILE HOME INSERT PAGE LAYOUT FORMULAS DATAs ? X A FILE - Sign In * s , B High-Low Method; Contribution Format Income Statement - Excel HOME INSERT PAGE LAYOUT FORMULAS
Answer #1

As a rule, only one (1) question can be answered at a time. In case no specific question is requested (maximum of 1), then the first questions is itself only answered.

In the case of JPL Inc.:

1. Compute the CM ratio and variable expense ratio
Selling price per unit (given) (SP) =B5 per unit 75 per unit
Variable expenses per unit (given) (VC) =B6 per unit 45 per unit
Contribution margin per unit [CM = (SP-VC)] =B12-B13 per unit 30 per unit
CM Ratio (CM / SP) =B14/B12 0.4
Variable expense ratio (VC / SP) =B13/B12 0.6
2. Compute the break-even point
Break-even in unit sales (BEU = Fixed expenses / CM) =B7/B14 units 7000 units
Break-even in sales (BE$ = BEU * SP) =B20*B5 525000
3. Compute Margin of safety
Margin of safety in dollars (MOS $ = Current sales - BE$) =(B4*B5)-B21 315000
Margin of safety in percentage (MOS $ / Current sales) =(B24/(B4*B5))*100 37.5
4. Compute degree of operating leverage
Sales =B4*B5 840000
Variable expenses =B4*B6 504000
Contribution Margin =B28-B29 336000
Fixed expense =B7 210000
Net operating income =B30-B31 126000
Degree of operating leverage =B30/B32 2.666667
(Contribution Margin / Net operating income)